# Must Read: A Systematic Survey of Computational Persuasion

### v1

@@ 4.2.1. Negotiation
Similarly, He et al. (2018) observed two agents that interact with each other to bargain on goods. They proposed an approach based on coarse dialogue acts, where they separate strategy and generation. They set the strategy using supervised learning, reinforcement learning, or domain-specific knowledge. This is used to select a dialogue act, then they generate a response. They show that this proposed method demonstrates more human-like negotiation and a higher task success rate based on human evaluations. Later works bring more focus to planning and persuasive strategies for negotiation. DialoGraph (Joshi et al., 2021) incorporated Graph Neural Networks to model sequences of strategies for negotiation, and incorporate them into a negotiation dialog system. (Chawla et al., 2021b) presented a dataset of negotiation dialogues and a multi-task framework for recognizing the persuasion strategies used in an utterance. Most recently, Bianchi et al. (2024) created NegotiationArena, a framework used to assess how well LLMs negotiate with each other to maximize their profits and resources. The results showed that GPT-4 is overall the best negotiating LLM. Their experiments highlight some interesting findings, such as cunning and desperate behaviors increase win rate and payoff.
The increasing persuasive capabilities of LLMs open new possibilities for negotiation, where models can serve as collaborative agents, mediators, or decision-support tools that guide discussions toward favorable and fair outcomes. However, these enhanced abilities also raise questions about how susceptible such systems may be when operating in multi-agent environments or interacting with humans who attempt to manipulate them. Understanding both the strengths and vulnerabilities of persuasive LLMs will be essential for building negotiation agents that are both effective and robust.
@@ 4.2.2. Debate
In contrast to the cooperative dynamics of negotiation, debate is inherently adversarial, with agents taking opposing positions and attempting to persuade an audience or judge of their stance. Michael et al. (2023) explored this setting by prompting two expert LLMs to debate answers to reading comprehension questions to persuade a non-expert human judge. The back-and-forth exchange exposes flaws and inconsistencies in each model’s reasoning, ultimately helping the human identify the correct answer. This work highlights the potential of debate as a supervision mechanism for aligning expert models and underscores the importance of studying persuasion as a core component of such interactions. Du et al. (2024) proposed a multi-agent debate framework in which LLMs iteratively propose and critique responses across several rounds before converging on a final answer. This debate-driven strategy was shown to improve mathematical reasoning, strategic thinking, and the factual correctness of generated content. Together, these studies demonstrate that debate not only facilitates persuasion but also contributes to model alignment and output reliability through adversarial reasoning.
@@ 4.2.3. Jailbreaking
Harmful/Toxic Content Generation. Jailbreaking LLMs means subverting their safety measures and producing harmful, biased, or sensitive generations. Recent work (Singh et al., 2023; Zeng et al., 2024) highlights a vulnerability in LLMs, where persuasion-based jailbreaks can bypass existing safety mechanisms. Unlike traditional algorithmic attacks (e.g., optimization-based attack), this study explores how human-like persuasion communication can manipulate LLMs to generate restricted content. Singh et al. (2023) demonstrates that LLMs can be deceived using tailored persuasion strategies, where prompts that mimic human-like persuasive communication are systematically crafted by leveraging authority, trust, and social proof. Zeng et al. (2024) develop a Persuasive Paraphraser which systematically rephrases harmful queries using a persuasion taxonomy grounded in social science, creating Persuasive Adversarial Prompts (PAPs). In a single-turn setting, they measure the Attack Success Rate (ASR), where PAPs achieve an ASR of over 92% on LLaMA-2, GPT-3.5, and GPT-4, surpassing conventional jailbreak methods. A key finding is that larger LLMs exhibit greater susceptibility to such persuasive attacks, likely due to their improved contextual reasoning and user engagement capabilities. They also evaluate existing defense mechanisms, revealing that existing mutation- and detection-based defenses are insufficient, while adaptive system-level defenses (e.g., reinforcement against persuasion) provide some mitigation, but remain limited. Moreover, Li et al. (2024) introduces multi-turn prompts with the persuasion technique, where users repeatedly apply persuasion techniques over conversations, leading LLMs to generate harmful content in multi-turn dialogues. These findings emphasize the need to rethink AI safety in the context of human-like communication, advocating for more robust defenses against social science techniques to prevent unintended model behavior.
@@ 5. Safeguarding Persuasion
As persuasive AI systems become increasingly influential, it becomes critical to ensure their responsible deployment. Safeguarding persuasion involves a wide range of challenges, including but not limited to understanding when and how persuasion should be employed, identifying and mitigating unwarranted or manipulative persuasive tactics, and calibrating the susceptibility of AI systems to persuasion without making them overly rigid. In this section, we review early work aimed at reducing the risks of unsafe persuasion and developing models that can more selectively decide when to accept, reject, or generate persuasive content. Compared to the other two areas covered in this survey—Evaluating Persuasion (§ 3) and Generating Persuasion (§ 4)—although equally important, this area has the least amount of prior work, in part because concerns around the safe use of persuasive AI have only recently begun to take shape with the advance of increasingly persuasive models.
@@ 5.2. Selective Acceptance of Persuasion
In addition to determining when and how to persuade, models must also develop the capacity to discern when to accept influence and when to resist it. We refer to this capability as selective acceptance of persuasion–the ability of a model to intelligently evaluate incoming persuasive attempts and decide whether to adopt or reject them based on context, alignment, and reliability. In § 2, we reviewed various strategies for generating persuasive content. Complementing this view, Dutt et al. (2021) introduced a novel line of work that shifts focus to resistance strategies, which are the methods individuals use to counter or reject persuasive attempts. Their framework, ResPer, drew from cognitive and social psychology to identify and categorize these strategies in real-time dialogue. Their approach revealed how resistance unfolds dynamically within conversations, exposing asymmetries in power and influence. Crucially, their findings showed that resistance strategies are often more predictive of conversation outcomes than persuasion tactics themselves. Such insights could inform the design of AI systems that are more discerning in responding to persuasive input, making them more receptive to beneficial influence while remaining resilient to manipulative attempts.
@@
While ResPer emphasized how models can push back, Stengel-Eskin et al. (2025) focused on when models should do so by proposing Persuasion-Balanced Training (PBT), a method that teaches models to selectively accept helpful (positive) persuasion while resisting harmful (negative) attempts. They simulate dialogues between persuader and persuadee agents, where the goal is to reach a consensus on a given question. These interactions are structured as recursive dialogue trees, from which preference data is derived. Using DPO loss (Rafailov et al., 2023), they train models to favor dialogue paths that lead to correct answers, regardless of whether persuasion was accepted or resisted. Their findings show that PBT significantly reduces misinformation and flipflopping, and improves collaboration in multi-agent debates, outperforming models trained only to resist or accept persuasion. Taken together, ResPer and PBT offer complementary perspectives on selective acceptance: the former provides a taxonomy of resistance strategies, while the latter offers a learning framework for context-sensitive influence evaluation. Both are essential steps toward developing models that are not only persuasive but also resilient and discerning.
@@
Still, real-world deployments pose additional challenges. In open-ended interactions with unfamiliar agents, intent is hard to infer, and deciding to accept or reject persuasion is inherently more nuanced. Sycophancy, the tendency to favor flattering or agreeable responses, presents an added challenge to learning selective persuasion. Both humans and preference models have been shown to favor convincingly written sycophantic responses over those that correct false beliefs in many cases (Sharma et al., 2024). This could lead to models falsely accepting harmful persuasion while still rejecting helpful ones if not handled with care. To build trust in persuasive systems, we must deepen our understanding of what models internalize during training and how they make influence-sensitive decisions.
To identify and address such vulnerabilities, red-teaming has emerged as a powerful tool. Scalable LLM-to-LLM red-teaming frameworks enable adversarial agents to systematically provoke undesirable behavior in target models before deployment (Perez et al., 2022). In persuasion-specific settings, red-team agents can manipulate models into accepting misleading influence, generating examples of where discernment fails. These adversarial conversations can then be used to train more robust models, helping them recognize patterns that lead to harmful persuasion and improve their ability to respond appropriately in real-world interactions.
@@ 6. Persuasion Beyond English Text
Mutimodal Persuasion. Although most computational research on persuasion has focused on textual data, human persuasion is inherently multimodal and often conveyed through speech, visuals, or a combination of modalities. Understanding how persuasive cues operate across these forms is crucial for developing more generalizable and context-aware persuasive systems. Early work by Park et al. (2014) explored multimodal persuasion by predicting persuasiveness in movie review videos using speech, text, and visual features. Building on this, Bai et al. (2023) proposed the Multimodal Persuasion Prediction (M2P2) framework, which models both the outcome and intensity of persuasion in debates. Later, Liu et al. (2022) introduced the ImageArg dataset, designed to support image-based persuasion detection. Other efforts have leveraged transcribed video data from social deduction games to study persuasive behaviors in complex, interactive settings (Lai et al., 2023). In line with textual strategy detection, researchers have also explored persuasive strategies in visual advertisements (Kumar et al., 2023), and in meme-based content through shared tasks organized at SemEval (Dimitrov et al., 2021, 2024). Xu et al. (2024b) demonstrated that vision-language models can be poisoned to generate misleading yet persuasive multimodal narratives. These efforts lay the groundwork for future research in persuasive content generation and safety mechanisms for multimodal persuasion.
Multilingual and Culture-Aware Persuasion. As with many NLP tasks, research on persuasion has predominantly focused on English, thanks to the availability of large-scale datasets and language resources. However, real-world persuasive communication spans a multitude of languages and cultural contexts, where data scarcity and linguistic diversity introduce additional challenges. The task of modeling persuasion becomes significantly more complex in multilingual and cross-cultural settings, hence, it remains relatively underexplored. One of the earliest efforts toward multilingual persuasion modeling is presented by Pöyhönen et al. (2022), who introduced a dataset of persuasion in video content across multiple languages. More recently, Piskorski et al. (2023) organized a SemEval shared task on persuasive strategy classification in a multilingual news corpus. Substantial research gaps remain, and fundamental tasks such as persuasion detection and strategy classification have yet to be fully extended to non-English languages. Furthermore, it remains an open question how cultural and linguistic context influences the choice and effectiveness of persuasive strategies. Addressing these gaps is essential for building persuasive systems that are culturally sensitive and globally applicable.
@@ 7.1. AI as Persuader
The generation of persuasive content using LLMs offers significant promise, but also presents serious risks. Although recent studies have shown that LLMs are becoming increasingly effective at persuasion (Durmus et al., 2024; OpenAI, 2024, 2025; Bozdag et al., 2025; Singh et al., 2024), current research has merely scratched the surface, leaving much to be explored in understanding the capabilities, limitations, and broader implications of persuasive LLMs.
@@ 7.1.2. Mechanics of Persuasion Generation
While LLMs are demonstrating notable improvements in persuasive capability (OpenAI, 2025, 2024; Durmus et al., 2024; Phuong et al., 2024), we still lack a clear understanding of what makes them effective (or ineffective) persuaders. Much of the underlying mechanics that drive LLMs’ persuasive success remains opaque, limiting our ability to reliably predict, enhance, or safeguard their behavior. Four major research gaps are:
@@
(2) Pro-social Applications of Persuasion: After early work such as Wang et al. (2019) which examined persuasion in donation settings, research has rarely explored how persuasive capabilities could be leveraged for socially beneficial domains. Fields such as education, healthcare, and personal well-being remain underexamined, despite their clear societal importance. Understanding how LLMs can adapt persuasive strategies to support learning, promote healthy behaviors, or enhance mental well-being presents a critical opportunity for future work. Moreover, studying persuasion in these contexts can inform the development of safeguards that prevent coercion and promote genuinely pro-social influence.
(3) Target-Specific and Adaptive Persuasion: While LLMs show some contextual adaptation, their ability to tailor persuasive strategies to individual users remains largely unexplored. Advancing this line of research will require controlled studies where models are given user background information and tasked with generating personalized persuasive messages, with effectiveness compared against generic outputs. Future work should also explore adaptive persuasion in multi-turn settings, where models refine strategies based on user feedback or environmental signals, potentially leveraging iterative reasoning frameworks such as ReAct (Yao et al., 2023). Framed as a preference learning and alignment problem, adaptive persuasion requires models to infer and respect user values, goals, and receptivity to generate effective and ethically grounded influence.
(4) Learning Persuasive Policies through Reinforcement: It remains an open question whether LLMs can be systematically trained to improve their persuasive decision-making. Current approaches based on prompt engineering or supervised fine-tuning do not equip models with dynamic, adaptive policies for persuasion. Future research should pursue persuasive policy learning, where models learn to generate persuasive content through reinforcement signals that reward both efficacy and ethical responsibility. This requires designing reward models that balance persuasive success with user well-being, fairness, and trustworthiness.

### v2

@@ 4.2.1. Negotiation
Similarly, He et al. (2018) observed two agents that interact with each other to bargain on goods. They proposed an approach based on coarse dialogue acts, where they separate strategy and generation. They set the strategy using supervised learning, reinforcement learning, or domain-specific knowledge. This is used to select a dialogue act, then they generate a response. They show that this proposed method demonstrates more human-like negotiation and a higher task success rate based on human evaluations. Later works bring more focus to planning and persuasive strategies for negotiation. DialoGraph (Joshi et al., 2021) incorporated Graph Neural Networks to model sequences of strategies for negotiation, and incorporate them into a negotiation dialog system. (Chawla et al., 2021a) presented a dataset of negotiation dialogues and a multi-task framework for recognizing the persuasion strategies used in an utterance. Most recently, Bianchi et al. (2024) created NegotiationArena, a framework used to assess how well LLMs negotiate with each other to maximize their profits and resources. The results showed that GPT-4 is overall the best negotiating LLM. Their experiments highlight some interesting findings, such as cunning and desperate behaviors increase win rate and payoff.
The increasing persuasive capabilities of LLMs open new possibilities for negotiation, where models can serve as collaborative agents, mediators, or decision-support tools that guide discussions toward favorable outcomes. However, these enhanced abilities also raise questions about how susceptible such systems may be when operating in multi-agent environments or interacting with humans who attempt to manipulate them. Understanding both the strengths and vulnerabilities of persuasive LLMs will be essential for building negotiation agents that are both effective and robust.
@@ 4.2.2. Debate
In contrast to the cooperative dynamics of negotiation, debate is inherently adversarial, with agents taking opposing positions and attempting to persuade an audience or judge of their stance. A foundational step toward computational debate was Slonim et al. (2021)’s Project Debater, an autonomous system capable of mining, organizing, and rebutting arguments in live debates against human experts, illustrating that large-scale argument generation and persuasion can be implemented in AI. In subsequent research, Michael et al. (2023) explored prompting two expert LLMs to debate answers to reading comprehension questions to persuade a non-expert human judge. The back-and-forth exchange exposes flaws and inconsistencies in each model’s reasoning, ultimately helping the human identify the correct answer. This work highlights the potential of debate as a supervision mechanism for aligning expert models and underscores the importance of studying persuasion as a core component of such interactions. Du et al. (2024) proposed a multi-agent debate framework in which LLMs iteratively propose and critique responses across several rounds before converging on a final answer. This debate-driven strategy was shown to improve mathematical reasoning, strategic thinking, and the factual correctness of generated content. Together, these studies demonstrate that debate not only facilitates persuasion but also contributes to model alignment and output reliability through adversarial reasoning.
@@ 4.2.3. Jailbreaking
Harmful/Toxic Content Generation. Jailbreaking LLMs means subverting their safety measures and producing harmful, biased, or sensitive generations. Recent work (Singh et al., 2023; Zeng et al., 2024) highlight a vulnerability in LLMs, where persuasion-based jailbreaks can bypass existing safety mechanisms. Singh et al. (2023) demonstrates that LLMs can be deceived using tailored persuasion strategies, where persuasive prompts are systematically crafted by leveraging authority, trust, and social proof. Zeng et al. (2024) develop a Persuasive Paraphraser which rephrases harmful queries using a persuasion taxonomy grounded in social science. A key finding is that larger LLMs exhibit greater susceptibility to such persuasive attacks, likely due to their improved contextual reasoning and user engagement capabilities. They also evaluate existing defense mechanisms, revealing that existing mutation- and detection-based defenses are insufficient, while adaptive system-level defenses (e.g., reinforcement against persuasion) provide some mitigation, but remain limited. Moreover, Li et al. (2024) introduces multi-turn prompts with the persuasion technique, where users repeatedly apply persuasion techniques over conversations, leading LLMs to generate harmful content in multi-turn dialogues. These findings emphasize the need to rethink AI safety in the context of human-like communication, advocating for more robust defenses against social science techniques to prevent unintended model behavior.
@@ 5. Safeguarding Persuasion
While persuasive communication can be beneficial, such as helping users make informed, healthy, or prosocial choices, it also carries risks of manipulation, bias, and erosion of autonomy. Responsible AI frameworks emphasize transparency, accountability, fairness, and non-maleficence (Radanliev, 2025; Radanliev et al., 2024). These principles extend naturally to persuasive models, which must balance efficacy with respect for user consent and epistemic integrity. A central issue in persuasion ethics is distinguishing virtuous vs. vicious persuasion. As argued by Godber and Origgi (2023), legitimate (virtuous) persuasion engages people’s deliberative and epistemic capacities, enabling reflective choice and preserving intellectual autonomy. Manipulative or propagandistic (vicious) persuasion instead exploits cognitive or emotional vulnerabilities to advance the persuader’s interests at the expense of the audience’s welfare. From a computational perspective, Bassi et al. (2024) observe that current NLP and ML systems seldom encode a principled distinction between ethical and unethical persuasion. Embedding such conceptual clarity into computational frameworks is essential for developing persuasion systems accountable to humanistic values.
As persuasive AI systems become increasingly influential, it becomes critical and urgent to ensure their responsible deployment. Safeguarding persuasion involves a wide range of challenges, including but not limited to understanding when and how persuasion should be employed, identifying and mitigating unwarranted or manipulative persuasive tactics, and calibrating the susceptibility of AI systems to persuasion without making them overly rigid. In this section, we review early work aimed at reducing the risks of unsafe persuasion and developing models that can more selectively decide when to accept, reject, or generate persuasive content. Compared to the other two areas covered in this survey—Evaluating Persuasion (§ 3) and Generating Persuasion (§ 4)—although equally important, this area has the least amount of prior work, in part because concerns around the safe use of persuasive AI have only recently begun to take shape with the advance of increasingly persuasive models.
@@ 5.2. Selective Acceptance of Persuasion
In addition to determining when and how to persuade, models must also develop the capacity to discern when to accept influence and when to resist it. We refer to this capability as selective acceptance of persuasion–the ability of a model to intelligently evaluate incoming persuasive attempts and decide whether to adopt or reject them based on context, alignment, and reliability. Dutt et al. (2021) introduced a novel line of work that shifts focus to resistance strategies, which are the methods individuals use to counter or reject persuasive attempts. Their framework, ResPer, drew from cognitive and social psychology to identify and categorize these strategies in real-time dialogue. Their approach revealed how resistance unfolds dynamically within conversations, exposing asymmetries in power and influence. Crucially, their findings showed that resistance strategies are often more predictive of conversation outcomes than persuasion tactics themselves. Such insights could inform the design of AI systems that are more discerning in responding to persuasive input, making them more receptive to beneficial influence while remaining resilient to manipulative attempts.
@@
Stengel-Eskin et al. (2025) focused on when models should do so by proposing Persuasion-Balanced Training (PBT), a method that teaches models to selectively accept helpful (positive) persuasion while resisting harmful (negative) attempts. They simulate dialogues between persuader and persuadee agents, where the goal is to reach a consensus on a given question. These interactions are structured as recursive dialogue trees, from which preference data is derived. Using DPO loss (Rafailov et al., 2023), they train models to favor dialogue paths that lead to correct answers, regardless of whether persuasion was accepted or resisted. Their findings show that PBT significantly reduces misinformation and flipflopping, and improves collaboration in multi-agent debates, outperforming models trained only to resist or accept persuasion. Taken together, ResPer and PBT offer complementary perspectives on selective acceptance: the former provides a taxonomy of resistance strategies, while the latter offers a learning framework for context-sensitive influence evaluation.
@@
Still, real-world deployments pose additional challenges. In open-ended interactions with unfamiliar agents, intent is hard to infer, and deciding to accept or reject persuasion is inherently more nuanced. Sycophancy, the tendency to favor flattering or agreeable responses, presents an added challenge to learning selective persuasion. Both humans and preference models have been shown to favor convincingly written sycophantic responses over those that correct false beliefs in many cases (Sharma et al., 2024). This could lead to models falsely accepting harmful persuasion while still rejecting helpful ones. To build trust in persuasive systems, we must deepen our understanding of what models internalize during training and how they make influence-sensitive decisions. Red-teaming has emerged as a powerful tool for identifying such vulnerabilities. Scalable LLM-to-LLM red-teaming frameworks enable adversarial agents to systematically provoke undesirable behavior in target models before deployment (Perez et al., 2022). In persuasion-specific settings, red-team agents can manipulate models into accepting misleading influence, generating examples of where discernment fails. These adversarial conversations can then be used to train more robust models, helping them recognize patterns that lead to harmful persuasion and improve their ability to respond appropriately in real-world interactions.
@@ 6. Persuasion Beyond English Text
Mutimodal Persuasion. Although most computational research on persuasion has focused on textual data, human persuasion is inherently multimodal and often conveyed through speech, visuals, or a combination of modalities. Understanding how persuasive cues operate across these forms is crucial for developing more generalizable and context-aware persuasive systems. Early work by Park et al. (2014) explored multimodal persuasion by predicting persuasiveness in movie review videos using speech, text, and visual features. Building on this, Bai et al. (2023) proposed the Multimodal Persuasion Prediction (M2P2) framework, which models both the outcome and intensity of persuasion in debates. Later, Liu et al. (2022) introduced the ImageArg dataset, designed to support image-based persuasion detection. Other efforts have leveraged transcribed video data from social deduction games to study persuasive behaviors in complex, interactive settings (Lai et al., 2023). Researchers have also explored persuasive strategies in visual advertisements (Kumar et al., 2023), and in meme-based content through shared tasks organized at SemEval (Dimitrov et al., 2021, 2024). Xu et al. (2024b) demonstrated that vision-language models can be poisoned to generate misleading yet persuasive multimodal narratives. These efforts lay the groundwork for future research in persuasive content generation and safety mechanisms for multimodal persuasion.
Multilingual and Culture-Aware Persuasion. As with many NLP tasks, research on persuasion has predominantly focused on English, thanks to the availability of large-scale datasets and language resources. However, real-world persuasive communication spans a multitude of languages and cultural contexts, where data scarcity and linguistic diversity introduce additional challenges. The task of modeling persuasion becomes significantly more complex in multilingual and cross-cultural settings, hence, it remains relatively underexplored. One of the earliest efforts toward multilingual persuasion modeling is presented by pöyhönen2022multilingualpersuasiondetectionvideo, who introduced a dataset of persuasion in video content across multiple languages. More recently, Piskorski et al. (2023) organized a SemEval shared task on persuasive strategy classification in a multilingual news corpus. Substantial research gaps remain, and fundamental tasks such as persuasion detection and strategy classification have yet to be fully extended to non-English languages. Furthermore, it remains an open question how cultural and linguistic context influences the choice and effectiveness of persuasive strategies. Addressing these gaps is essential for building persuasive systems that are culturally sensitive and globally applicable.
@@ 7.1. AI as Persuader
The generation of persuasive content using LLMs offers significant promise, but also presents serious risks. Although recent studies have shown that LLMs are becoming increasingly effective at persuasion (Durmus et al., 2024; OpenAI, 2024b, 2025; Bozdag et al., 2025; Singh et al., 2024), current research has merely scratched the surface, leaving much to be explored in understanding the capabilities, limitations, and broader implications of persuasive LLMs.
@@ 7.1.2. Mechanics of Persuasion Generation
While LLMs are demonstrating notable improvements in persuasive capability (OpenAI, 2025, 2024b; Durmus et al., 2024; Phuong et al., 2024), we still lack a clear understanding of what makes them effective (or ineffective) persuaders. Much of the underlying mechanics that drive LLMs’ persuasive success remains opaque, limiting our ability to reliably predict, enhance, or safeguard their behavior. Four major research gaps are:
@@
(2) Pro-social Applications of Persuasion: After early work such as Wang et al. (2019) which examined persuasion for donation, research has rarely explored how persuasive capabilities could be leveraged for socially beneficial domains. Fields such as education, healthcare, and personal well-being remain underexamined, despite their clear societal importance. Understanding how LLMs can adapt persuasive strategies to support learning, promote healthy behaviors, or enhance mental well-being presents a critical opportunity for future work. Moreover, studying persuasion in these contexts can inform the development of safeguards that prevent coercion and promote genuinely pro-social influence.
(3) Target-Specific and Adaptive Persuasion: While LLMs show some contextual adaptation, their ability to tailor persuasive strategies to individual users remains largely unexplored. Advancing this line of research will require controlled studies where models are given user background information and tasked with generating personalized persuasive messages, with effectiveness compared against generic outputs. Future work should also explore adaptive persuasion in multi-turn settings, where models refine strategies based on user feedback or environmental signals. Earlier decision-theoretic approaches have treated persuasion as an interaction between agents with separate objectives, where user preferences are learned to guide adaptive argument choices (Donadello et al., 2022). Framed as a preference learning and alignment problem, adaptive persuasion requires models to infer and respect user values, goals, and receptivity to generate effective and ethically grounded influence.
(4) Learning Persuasive Policies through Reinforcement: It remains an open question whether LLMs can be systematically trained to improve their persuasive decision-making. Current approaches based on prompt engineering or supervised fine-tuning do not equip models with dynamic, adaptive policies for persuasion. Future research should pursue persuasive policy learning, where a policy represents the model’s learned mapping from dialogue states to persuasive actions that optimize a defined objective—be it ethical influence, pro-social encouragement, commercial engagement, or adversarial robustness. In this context, models learn to generate persuasive content through reinforcement signals that reward both efficacy and ethical responsibility. This requires designing reward models that balance persuasive success with user well-being, fairness, and trustworthiness.